OpenStudy (anonymous):

If a^p = a^3t and a^t= a^5pr and p does not equal 0, then a= ?

OpenStudy (anonymous):

answer choices: 1 2 3 4 5

OpenStudy (primeralph):

Is this correct? a^5pr or is it a^5p ?

OpenStudy (anonymous):

Is this \(a^p=a^{3t}\) and \(a^t=a^{5p}\) ... ?

OpenStudy (anonymous):

\(a\) could be either \(0,1\)

OpenStudy (anonymous):

it is a^t=a^(5pr)

OpenStudy (anonymous):

Assume \(a\ne 0,1\) then we may equate exponents \(p=3t,t=5p\implies t=15t\). \(t=0\) therefore and so does \(p\) -- we've reached a contradiction. Therefore \(a=0\) or \(a=1\)

OpenStudy (anonymous):

what the hell is \(r\)?

OpenStudy (anonymous):

are there any conditions on \(a,r\)?

OpenStudy (anonymous):

If a^p = a^(3t) and a^t= a^(5pr) and p does not equal 0, then what is the value of r? @oldrin.bataku @primeralph My bad!

OpenStudy (anonymous):

Are we told \(a\ne0,1\)? otherwise it could be literally anything

OpenStudy (anonymous):

well, any real number

OpenStudy (anonymous):

well, it says p doesn't equal zero, but that's it.

OpenStudy (anonymous):

If this is supposed to hold for any \(a\) I suppose we could argue \(r=1/15\) is a solution... see:$$a^p=a^{3t}\implies p=3t\\a^t=a^{5pr}\implies t=5pr\\t=5pr=3(5t)r=15tr\\\frac1{15}=r$$